Find the parametric equation of the line that passes through the points (0,1,1) and (1,0,-1).


1
Expert's answer
2020-07-07T16:46:01-0400

Soiution:


"\begin{cases}
 x=x_1+(x_2-x_1)\times\lambda \\
 x=y_1+(y_2-y_1)\times\lambda \\ z=z_1+(z_2-z_1)\times\lambda \\
 \end{cases} \Rightarrow\\ \Rightarrow \begin{cases}
 x=0+(1-0)\times\lambda \\
 у=1+(0-1)\times\lambda \\ z=1+(-1-1)\times\lambda \\
 \end{cases} \Rightarrow \\ \Rightarrow \begin{cases}
 x=\lambda \\
 у=1-\lambda \\ z=1-2\times\lambda \\
 \end{cases} \\Answer: \\\;\;\;\;\; \begin{cases}
 x=\lambda \\
 у=1-\lambda \\ z=1-2\times\lambda \\
 \end{cases}"